# FedALT: Federated Fine-Tuning through Adaptive Local Training with Rest-of-World LoRA

### v2

@@ 4.3 Sensitivity Analysis
Table 4: Performance under varying heterogeneity levels. Evaluation on Bloom-560M and LLaMA2-7B models across three levels of task heterogeneity. See section 4.3.
Methods Bloom-560M LLaMA2-7B Low Mild High Low Mild High Local 48.65 50.21 44.40 60.59 66.67 62.86 FedIT 52.15 49.03 44.33 62.24 64.53 62.19 FFA-LoRA 51.33 47.32 40.44 61.07 62.15 56.78 FedSA 53.03 52.55 46.98 62.87 67.44 63.47 FedDPA 53.12 52.06 46.99 62.06 67.49 64.64 PF2LoRA 51.10 51.97 45.87 63.56 67.60 64.44 FDLoRA 52.51 51.42 45.54 62.42 66.56 65.17 FedALT 54.48 54.89 48.09 64.66 70.48 67.55
@@
Effect of Local Epochs. We also investigate the effect of varying local training epochs (5 and 10), while keeping all other configurations unchanged. As presented in fig. 4(c), our method outperforms all baselines under both settings, further highlighting its robustness to different local training epochs.
(a) Effect of number of clients. See section 4.3.
(b) Effect of LoRA rank. See section 4.3.
(c) Effect of local epochs. See section 4.3.
5 Related Work
FedAvg [27] addresses privacy and communication efficiency in federated learning by aggregating locally trained model parameters to construct a shared global model. With the growing interest in LLMs, FL has emerged as a promising framework for fine-tuning LLMs while preserving data privacy. Recent works have explored federated fine-tuning of LLMs by either updating the entire model [20] or applying PEFT techniques such as prefix-tuning [33] and LoRA [49]. Subsequent studies [1, 5, 43, 3, 34] have tackled key challenges in this domain, including heterogeneous LoRA configurations, initialization strategies, and server-side aggregation bias. Personalized federated fine-tuning methods [15, 45, 30] have proposed combining global and local LoRA modules, but often suffer from cross-client interference during aggregation or depend on server-side data. Our method introduces a robust and scalable approach to personalized fine-tuning across heterogeneous clients. A more comprehensive review of related work is provided in appendix C.
6 Conclusion
@@ Appendix A Datasets
In this paper, we follow the setup proposed in [45] to construct federated datasets for heterogeneous client settings using the Flan collection [6]. The Flan benchmarks comprises over 12 NLP task categories, each containing multiple datasets. These tasks, derived from diverse contextual domains, inherently involve complex distribution shifts. We consider three federated dataset configurations. In the High Heterogeneous setting, 8 clients are assigned entirely different NLP tasks, resulting in a total of 8 distinct task types. In the Low Heterogeneous setting, 8 clients are distributed across only two tasks—Sentiment Analysis and Natural Language Inference—with all clients working on the same task drawing data from a common source, thereby minimizing heterogeneity. The Mild Heterogeneous setting assigns 8 clients to datasets spanning 6 different NLP tasks. To better simulate realistic FL environments, where each client has limited access to large-scale data, we apply a downsampling strategy: each client is assigned 600 training samples and 300 testing samples. For our extended evaluations with 16 and 24 clients, we maintain the same total dataset size while distributing it across more clients. Specifically, in the 16-client configuration, each client receives 300 training samples and 150 testing samples, while in the 24-client setting, each client is allocated 200 training samples and 100 testing samples. The specific tasks and datasets used in each configuration are provided in tables 5, 6 and 7. While some datasets appear in multiple configurations, the sample selections are made independently and randomly.
The original Flan tasks exhibit highly diverse formats, making standardized preprocessing and evaluation across LLMs challenging. Even within a single dataset (e.g., Sentiment140), sample formats can vary significantly, as illustrated in tables 8 and 9. To ensure consistency, we adopt a unified prompt template based on [45], shown in table 10. For evaluation, we follow FedDPA [45] and use ROUGE-1 as the metric across all tasks, providing a unified measure of model performance. This choice is particularly appropriate as Flan inherently structures NLP tasks in a generative format, where ROUGE-1 effectively captures lexical overlap between generated outputs and reference answers. Using ROUGE-1 enables direct comparability across diverse tasks and clients, which is essential for analyzing federated learning dynamics. While task-specific metrics might offer more nuanced evaluation for individual tasks, they would introduce incomparability issues that complicate assessment of the overall federated learning performance, making ROUGE-1 a methodologically sound choice.
@@
Template
Prompt Input Instruction: {instruction} Response:
@@ Appendix B Training Implementations
We implement all methods using PyTorch and conduct all experiments on six NVIDIA A100 GPUs. For the federated methods, we set the local epochs to 5 and the global rounds to 20. For the Local Only baseline, the total number of training rounds is set to 50. The learning rate is explored in the range $\{3\times 10^{-2},1\times 10^{-3},3\times 10^{-3},1\times 10^{-4},3\times 10^{-4},1\times 10^{-5}\}$, and the AdamW optimizer is utilized, following [45].
@@
Appendix C Related Work and Compared Methods
C.1 Parameter-Efficient Fine-Tuning
As the size of LLMs continues to grow, full parameter fine-tuning has become increasingly computationally and storage-intensive [13]. To address this issue, Parameter-Efficient Fine-Tuning (PEFT) methods [9, 11, 14, 25, 24, 23] have been developed to significantly reduce the number of trainable parameters. PEFT techniques introduce a limited set of additional trainable parameters to enhance model performance while keeping the majority of pre-trained parameters frozen. Among all existing PEFT methods, the most prominent approach is LoRA [16], which employs low-rank matrices to approximate the pre-trained weight matrix, updating only the low-rank components.
LoRA has gained significant attention recently, becoming a standard method for adapting LLMs like Llama under limited computational resources [22]. Several works have been proposed to improve LoRA [36, 22, 41, 32, 26]. For instance, a recent development, Hydra-LoRA [36], employs an asymmetric structure with a shared matrix for all samples and distinct matrices for each intrinsic component, thereby enhancing domain adaptation. While these advancements primarily focus on centralized learning settings, our work extends the application of PEFT methods to federated LLM fine-tuning, addressing the unique challenges posed by distributed and heterogeneous client data.
C.2 Federated Learning
The foundational work in federated learning, FedAvg [27], addresses privacy and communication efficiency by aggregating local model parameters to train a shared global model. Since then, numerous studies [39, 42, 46, 2] have focused on tackling various challenges within FL settings. One of the key challenges is data heterogeneity, which makes it difficult to train a single shared global model that performs well across all clients. To address this, Personalized Federated Learning(PFL) [35, 8, 7, 10] has emerged as an approach to adapt the global model to the specific needs of each client in the federation. Most PFL research has focused on addressing statistical differences in data distributions (e.g., label distribution skew) and has been applied primarily to smaller, standard models. In contrast, our work addresses the unique challenges posed by heterogeneous clients with diverse NLP tasks by leveraging LLM fine-tuning within the FL framework.
C.3 Federated Fine-Tuning
Several studies [1, 5, 43, 20, 44, 3] have explored federated fine-tuning approaches. [20] proposed federated fine-tuning with all parameters updated, while [33] introduced federated fine-tuning with PEFT using prefix-tuning. [49] was the first study to apply LoRA in a federated context. [1, 5] focused on federated fine-tuning where clients have different LoRA ranks, while [43] proposed strategies for LoRA initialization to achieve better performance. Another studies [3, 34] addressed server aggregation bias in LoRA-based federated fine-tuning.
The most relevant works to ours focus on personalized federated fine-tuning with LoRA [15, 45, 30]. FedDPA [45] introduced global and local LoRA components but sufferd from interference during global LoRA training and struggled to balance the two phases effectively. PF2LoRA [15] employed a bilevel framework to combine a shared adapter with client-specific adapters but faced interference issues during the aggregation of the shared adapter. FDLoRA [30] utilized personalized LoRA modules to initialize federated training, combining them with a global LoRA module via adaptive fusion; however, it relied on server-side datasets and was prone to interference during global LoRA aggregation, limiting its effectiveness in heterogeneous settings.
@@ C.3 Federated Fine-Tuning
FedIT [49]. FedIT integrates LoRA with the FedAvg framework, making it one of the earliest approaches to combine these techniques. However, under heterogeneous client data settings, FedIT suffers from harmful cross-client interference during the aggregation step, as highlighted in the motivational experiments. This interference often reduces its efficiency and, in some cases, results in worse performance than the Local Only method.
@@
FFA-LoRA [34]. FFA-LoRA aims to reduce communication costs by fixing the randomly initialized non-zero $\mathbf{A}$ matrices and fine-tuning only the zero-initialized $\mathbf{B}$ matrices. While this approach lowers communication overhead, it impairs the learning capacity of LoRA, as fixed $\mathbf{A}$ matrices limit the model’s ability to adapt to diverse client data, resulting in suboptimal performance.
@@
FedSA [12]. FedSA introduces an asymmetric approach to LoRA matrix updates in federated learning, where only the $\mathbf{A}$ matrices are shared with the server for aggregation, while the $\mathbf{B}$ matrices remain local to capture client-specific knowledge. However, even the client-specific $\mathbf{B}$ matrices may contain valuable knowledge that could benefit other clients. The failure to leverage this knowledge effectively results in reduced overall performance.
@@
PF2LoRA [15]. PF2LoRA adopts a two-level low-rank adaptation framework, where a common adapter shared across all clients is combined with client-specific lightweight adapters for personalization. This bilevel optimization framework aims to balance shared knowledge and personalization. However, the aggregation of the shared adapter during the server aggregation step can reintroduce harmful interference, limiting the method’s effectiveness in highly heterogeneous settings.
@@
FDLoRA [30]. FDLoRA follows a two-step process where personalized LoRA modules are first trained locally and then used to initialize federated training to obtain a global LoRA module. An adaptive fusion mechanism combines the two LoRA modules during fine-tuning. While this method improves personalization, it requires additional server-side datasets for effective global LoRA training. Moreover, the aggregation step during global LoRA training remains prone to interference, reducing its robustness in handling client heterogeneity.
@@
FedDPA [45]. FedDPA is the most closely related to our proposed method, as it also targets settings where clients possess heterogeneous tasks (not just label non-IID data). FedDPA employs both global and local LoRA modules, and enhances personalization by introducing a local fine-tuning step excluded from aggregation. However, our proposed FedALT departs from FedDPA in several critical ways: First, in both variants of FedDPA, the global LoRA module is trained independently of the local LoRA modules. During global training, the local modules are excluded, leading to persistent interference—similar to FedIT—and the local LoRA modules are never shared across clients. In contrast, FedALT introduces the Rest-of-World (RoW) LoRA, which captures knowledge from other clients via aggregation, allowing more effective information sharing. Second, FedALT employs a dynamic mixer that enables each client to adaptively balance the contributions of its Individual and RoW LoRA modules on a per-input basis, enabling finer-grained personalization. Third, FedDPA separates local and global LoRA training into distinct phases, introducing additional hyperparameters to tune the duration of each phase. Short local training reduces personalization; overly long training approximates local-only fine-tuning, diminishing the benefits of FL. In contrast, FedALT removes the need for phase separation and additional hyperparameters, simplifying training and improving robustness.
Appendix D The Algorithm of FedALT
@@
Input: Pretrained model $W_{0}$, number of clients $K$, communication rounds $T$
1 for each client $k$ do
2       Initialize Individual LoRA $(A_{k}^{L},B_{k}^{L})$ and Mixer $G_{k}$ randomly;
3       Initialize RoW LoRA $(A_{k}^{R},B_{k}^{R})\leftarrow 0$;
4
5for each round $t=1$ to $T$ do
       // Client-side training
6       for each client $k$ in parallel do
7             Receive RoW LoRA $(A_{k}^{R},B_{k}^{R})$ from server;
8             for each local epoch do
9                   Update $(A_{k}^{L},B_{k}^{L})$ and Mixer $G_{k}$ using gradient descent;
                   // Model output:
10                   $y=W_{0}x+\alpha_{k}(x)B_{k}^{L}A_{k}^{L}x+(1-\alpha_{k}(x))B_{k}^{R}A_{k}^{R}x$;
11                   where $\alpha_{k}(x)=\text{softmax}(G_{k}x)$
12            Upload $(A_{k}^{L},B_{k}^{L})$ to server;
13
14
      // Server-side aggregation
15       for each client $k$ do
16             Compute RoW LoRA: $A_{k}^{R}\leftarrow\frac{1}{K-1}\sum_{m\neq k}A_{m}^{L}$, $B_{k}^{R}\leftarrow\frac{1}{K-1}\sum_{m\neq k}B_{m}^{L}$;
17             Send $(A_{k}^{R},B_{k}^{R})$ to client $k$;
18
19
@@ 17             Send $(A_{k}^{R},B_{k}^{R})$ to client $k$;
Appendix E Additional Experiment Results
E.1 Training Workload
@@
Methods Trainable Parameters Avg Performance
RoW-Update (rank=8) 0.1283% 65.75
Fixed Weight (rank=16) 0.1244% 64.59
FedALT 0.0661% 67.55
E.2 Per-Client Performance
Due to space limitations, we reported only the average results for the Low Heterogeneous and Mild Heterogeneous settings in the main paper. Here, we present the detailed per-client performance in tables 13, 14, 15 and 16.
@@ E.2 Per-Client Performance
FedALT 54.06 49.18 53.34 55.78 42.94 45.51 67.57 67.49 54.48
Appendix F Convergence Analysis
@@ F.1 Notation and Setup
F.2 Assumptions
Assumption F.1 (Smoothness).
@@ F.2 Assumptions
Assumption F.2 (Strong Convexity).
@@
Assumption F.3 (Bounded RoW Sensitivity).
@@
$\|Z_{k}^{*}(\mathbf{A}_{k}^{R,1},\mathbf{B}_{k}^{R,1})-Z_{k}^{*}(\mathbf{A}_{k}^{R,2},\mathbf{B}_{k}^{R,2})\|\leq\beta\|(\mathbf{A}_{k}^{R,1},\mathbf{B}_{k}^{R,1})-(\mathbf{A}_{k}^{R,2},\mathbf{B}_{k}^{R,2})\|$ (9)
@@
F.3 Convergence Theorem
Theorem F.4 (Convergence of FedALT).
@@ F.3 Convergence Theorem
$L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ $\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})+\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{t,s+1}-Z_{k}^{t,s}\rangle$
@@
$L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ $\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
@@
$\quad+\frac{\mu}{2}\|-\eta\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (12)
@@
$\quad+\frac{\mu\eta^{2}}{2}\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (13)
$=L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-\eta(1-\frac{\mu\eta}{2})\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (14)
@@
$\geq$ $L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})+\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{*}-Z_{k}^{t,s}\rangle+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (15)
@@
$\geq$ $L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (16)
@@
$\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|\cdot\|Z_{k}^{t,s}-Z_{k}^{*}\|\geq\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{t,s}-Z_{k}^{*}\rangle$ (17)
@@
$\geq$ $L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (18)
@@
$\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}\geq 2\lambda(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (19)
@@
$\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-\eta(1-\frac{\mu\eta}{2})\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (20)
@@
$-2\eta(1-\frac{\mu\eta}{2})\lambda(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (21)
$=(1-2\eta\lambda(1-\frac{\mu\eta}{2}))(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})),$ (22)

### v3

@@ Introduction
Effect of Local Epochs. We investigate the effect of varying local training epochs (5 and 10), while keeping all other configurations unchanged. As presented in fig. 4(c), our method outperforms all baselines, further highlighting its robustness to different local training epochs.
Additional experiments on the effect of data heterogeneity can be found in the supplementary material.
Conclusion
@@
Acknowledgments
The work of Jieming Bian, Lei Wang and Jie Xu is partially supported by NSF under grants 2433886, 2505381 and 2515982. The work of Letian Zhang is partially supported by NSF under grant 2348279 and also supported by MTSU Stark Land project.
@@ Appendix A Datasets
In this paper, we follow the setup proposed in (yang2024dual) to construct federated datasets for heterogeneous client settings using the Flan collection (chung2024scaling). The Flan benchmarks comprise over 12 NLP task categories, each containing multiple datasets. These tasks, derived from diverse contextual domains, inherently involve complex distribution shifts. We consider three federated dataset configurations. In the High Heterogeneous setting, 8 clients are assigned entirely different NLP tasks, resulting in a total of 8 distinct task types. In the Low Heterogeneous setting, 8 clients are distributed across only two tasks—Sentiment Analysis and Natural Language Inference—with all clients working on the same task drawing data from a common source, thereby minimizing heterogeneity. The Mild Heterogeneous setting assigns 8 clients to datasets spanning 6 different NLP tasks. To better simulate realistic FL environments, where each client has limited access to large-scale data, we apply a downsampling strategy: each client is assigned 600 training samples and 300 testing samples. For our extended evaluations with 16 and 24 clients, we maintain the same total dataset size while distributing it across more clients. Specifically, in the 16-client configuration, each client receives 300 training samples and 150 testing samples, while in the 24-client setting, each client is allocated 200 training samples and 100 testing samples. The specific tasks and datasets used in each configuration are provided in Tables 3, 4, 5. Although some datasets appear in multiple configurations, sample selections are made independently and randomly.
The original Flan tasks exhibit highly diverse formats, making standardized preprocessing and evaluation across LLMs challenging. Even within a single dataset (e.g., Sentiment140), sample formats can vary significantly, as illustrated in tables 6 and 7. To ensure consistency, we adopt a unified prompt template based on (yang2024dual), shown in table 8. For evaluation, we follow FedDPA (yang2024dual) and use ROUGE-1 as the metric across all tasks, providing a unified measure of model performance. This choice is particularly appropriate as Flan inherently structures NLP tasks in a generative format, where ROUGE-1 effectively captures lexical overlap between generated outputs and reference answers. Using ROUGE-1 enables direct comparability across diverse tasks and clients, which is essential for analyzing federated learning dynamics. While task-specific metrics might offer more nuanced evaluation for individual tasks, they would introduce incomparability issues that complicate assessment of the overall federated learning performance, making ROUGE-1 a methodologically sound choice.
@@
Template
Prompt Input
Instruction: {instruction}
Response:
@@ Appendix B Training Implementations
We implement all methods using PyTorch and conduct all experiments on six NVIDIA A100 GPUs. For the federated methods, we set the local epochs to 5 and the global rounds to 20. For the Local Only baseline, the total number of training rounds is set to 50. The learning rate is explored in the range $\{3\times 10^{-2},1\times 10^{-3},3\times 10^{-3},1\times 10^{-4},3\times 10^{-4},1\times 10^{-5}\}$, and the AdamW optimizer is utilized, following (yang2024dual).
@@
Appendix C Compared Methods
@@
FedIT (zhang2024towards). FedIT integrates LoRA with the FedAvg framework, making it one of the earliest approaches to combine these techniques. However, under heterogeneous client data settings, FedIT suffers from harmful cross-client interference during the aggregation step, as highlighted in the motivational experiments. This interference often reduces its efficiency and, in some cases, results in worse performance than the Local Only method.
@@
FFA-LoRA (sun2024improving). FFA-LoRA aims to reduce communication costs by fixing the randomly initialized non-zero $\mathbf{A}$ matrices and fine-tuning only the zero-initialized $\mathbf{B}$ matrices. While this approach lowers communication overhead, it impairs the learning capacity of LoRA, as fixed $\mathbf{A}$ matrices limit the model’s ability to adapt to diverse client data, resulting in suboptimal performance.
@@
FedSA (guo2024selective). FedSA introduces an asymmetric approach to LoRA matrix updates in federated learning, where only the $\mathbf{A}$ matrices are shared with the server for aggregation, while the $\mathbf{B}$ matrices remain local to capture client-specific knowledge. However, even the client-specific $\mathbf{B}$ matrices may contain valuable knowledge that could benefit other clients. The failure to leverage this knowledge effectively results in reduced overall performance.
@@
PF2LoRA (hao2025personalized). PF2LoRA adopts a two-level low-rank adaptation framework, where a common adapter shared across all clients is combined with client-specific lightweight adapters for personalization. This bilevel optimization framework aims to balance shared knowledge and personalization. However, the aggregation of the shared adapter during the server aggregation step can reintroduce harmful interference, limiting the method’s effectiveness in highly heterogeneous settings.
@@
FDLoRA (qi2024fdlora). FDLoRA follows a two-step process where personalized LoRA modules are first trained locally and then used to initialize federated training to obtain a global LoRA module. An adaptive fusion mechanism combines the two LoRA modules during fine-tuning. While this method improves personalization, it requires additional server-side datasets for effective global LoRA training. Moreover, the aggregation step during global LoRA training remains prone to interference, reducing its robustness in handling client heterogeneity.
@@
FedDPA (yang2024dual). FedDPA is the most closely related to our proposed method, as it also targets settings where clients possess heterogeneous tasks (not just label non-IID data). FedDPA employs both global and local LoRA modules, and enhances personalization by introducing a local fine-tuning step excluded from aggregation. However, our proposed FedALT departs from FedDPA in several critical ways: First, in both variants of FedDPA, the global LoRA module is trained independently of the local LoRA modules. During global training, the local modules are excluded, leading to persistent interference—similar to FedIT—and the local LoRA modules are never shared across clients. In contrast, FedALT introduces the Rest-of-World (RoW) LoRA, which captures knowledge from other clients via aggregation, allowing more effective information sharing. Second, FedALT employs a dynamic mixer that enables each client to adaptively balance the contributions of its Individual and RoW LoRA modules on a per-input basis, enabling finer-grained personalization. Third, FedDPA separates local and global LoRA training into distinct phases, introducing additional hyperparameters to tune the duration of each phase. Short local training reduces personalization; overly long training approximates local-only fine-tuning, diminishing the benefits of FL. In contrast, FedALT removes the need for phase separation and additional hyperparameters, simplifying training and improving robustness.
Appendix D Discussion: Why a Single RoW LoRA?
Appendix E The Algorithm of FedALT
@@
0: Pretrained model $W_{0}$, number of clients $K$, communication rounds $T$
1:for each client $k$ do
2:  Initialize LoRA $(A_{k}^{L},B_{k}^{L})$ and mixer $G_{k}$ randomly
3:  Initialize RoW LoRA $(A_{k}^{R},B_{k}^{R})\leftarrow 0$
4:end for
5:for $t=1$ to $T$ do
6:  Client-side training:
7:  for each client $k$ in parallel do
8:   Receive RoW LoRA $(A_{k}^{R},B_{k}^{R})$ from server
9:   for each local epoch do
10:    Update $(A_{k}^{L},B_{k}^{L})$ and $G_{k}$ using gradient descent
11:    Compute output:
12:    $y=W_{0}x+\alpha_{k}(x)B_{k}^{L}A_{k}^{L}x+(1-\alpha_{k}(x))B_{k}^{R}A_{k}^{R}x$
13:    where $\alpha_{k}(x)=\text{softmax}(G_{k}x)$
14:   end for
15:   Upload $(A_{k}^{L},B_{k}^{L})$ to server
16:  end for
17:  Server-side aggregation:
18:  for each client $k$ do
19:   $A_{k}^{R}\leftarrow\frac{1}{K-1}\sum_{m\neq k}A_{m}^{L}$
20:   $B_{k}^{R}\leftarrow\frac{1}{K-1}\sum_{m\neq k}B_{m}^{L}$
21:   Send $(A_{k}^{R},B_{k}^{R})$ to client $k$
22:  end for
23:end for
Appendix F Additional Experiment Results
Training Workload
@@ Appendix F Additional Experiment Results
Methods Trainable Parameters Avg Performance RoW-Update (rank=8) 0.1283% 65.75 Fixed Weight (rank=16) 0.1244% 64.59 FedALT 0.0661% 67.55
@@
Effect of Data Heterogeneity
To further evaluate the robustness of our method under varying degrees of client heterogeneity, we design experiments across three settings. Our main experimental configuration, where each client is assigned a unique task, is denoted as the High Heterogeneous setting. We also include two alternative configurations. In the Low Heterogeneous setting, 8 clients are divided between only two tasks—Sentiment Analysis and Natural Language Inference—with clients sharing a task drawing data from the same source, resulting in minimal heterogeneity. The Mild Heterogeneous setting involves 8 clients assigned datasets spanning a total of 6 distinct NLP tasks. As shown in tables 11, 12, 13 and 14., FedALT consistently outperforms all baseline methods across all levels of heterogeneity, using both Bloom and LLaMA2 as pre-trained models. These results highlight the robustness and adaptability of FedALT in handling diverse and heterogeneous data distributions.
@@
[…]
Appendix G Convergence Analysis
@@
Assumptions
Assumption 1 (Smoothness).
@@
Assumption 2 (Strong Convexity).
@@
Assumption 3 (Bounded RoW Sensitivity).
@@
$\|Z_{k}^{*}(\mathbf{A}_{k}^{R,1},\mathbf{B}_{k}^{R,1})-Z_{k}^{*}(\mathbf{A}_{k}^{R,2},\mathbf{B}_{k}^{R,2})\|$
$\leq\beta\|(\mathbf{A}_{k}^{R,1},\mathbf{B}_{k}^{R,1})-(\mathbf{A}_{k}^{R,2},\mathbf{B}_{k}^{R,2})\|$ (9)
@@
Convergence Theorem
Theorem 1 (Convergence of FedALT).
@@
$L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
$\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
$\quad+\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{t,s+1}-Z_{k}^{t,s}\rangle$
@@
$L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
$\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
@@
$\quad+\frac{\mu}{2}\|-\eta\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$
@@
$\quad+\frac{\mu\eta^{2}}{2}\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$
$=L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
$\quad-\eta(1-\frac{\mu\eta}{2})\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (12)
@@
$\geq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
$\quad+\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{*}-Z_{k}^{t,s}\rangle$
$\quad+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (13)
@@
$\geq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
$\quad+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (14)
@@
$\geq\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{t,s}-Z_{k}^{*}\rangle$ (15)
@@
$\geq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
$\quad+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (16)
@@
$\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$
$\geq 2\lambda(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (17)
@@
$\quad-\eta(1-\frac{\mu\eta}{2})\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (18)
$\quad-2\eta(1-\frac{\mu\eta}{2})\lambda(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
$\quad-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (19)
$=(1-2\eta\lambda(1-\frac{\mu\eta}{2}))(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
$\quad-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})),$ (20)